OpenStudy (anonymous):

Is this a function???? (3,7), (3,8), (3,-2), (3,4), (3,1)

OpenStudy (fettiwap1738):

yes

OpenStudy (alekos):

I disagree

OpenStudy (anonymous):

ya its not i just got the anser

OpenStudy (anonymous):

thx for helping tho :)

OpenStudy (alekos):

no probs

OpenStudy (anonymous):

can u help me with one more question?

OpenStudy (alekos):

yep

OpenStudy (anonymous):

What is the common difference of the following arithmetic sequence? 13, -7, -27, -47, …

OpenStudy (alekos):

well. just subtract the second term with the first term

OpenStudy (anonymous):

do what?

OpenStudy (alekos):

If you have a arithmetic sequence a1,a2,a3,a4 and so on the difference = a2-a1

OpenStudy (anonymous):

ohhh ok so -13-7=-20

OpenStudy (anonymous):

ok thanks :)

OpenStudy (alekos):

no, it's actually -7-13 = -20 2nd term minus 1st term

OpenStudy (anonymous):

oh yea :P

OpenStudy (anonymous):

ok and how do i tell if a sequence is arithmetic?

OpenStudy (alekos):

all the terms have a common difference, d

OpenStudy (anonymous):

is this sequence arithmetic? 1, -7, -14, -21, …

OpenStudy (anonymous):

@alekos

OpenStudy (anonymous):

i think it itsn't

OpenStudy (alekos):

find a2-a1 and a3-a2